For all integers n and m, if n − m is even then n^3 − m^3

is even.


1
Expert's answer
2021-11-30T06:15:02-0500

Let "n" and "m" be integers such that "n\u2212m" is even. Therefore, "n\u2212m = 2k" for some "k ∈ \Z." Note that


"n^3-m^3=(n-m)(n^2+nm+m^2)"

"=2k(n^2+nm+m^2)=2(k(n^2+nm+m^2))"

Let "k(n^2+nm+m^2)." Since "k(n^2+nm+m^2)" is an integer, "n ^3 \u2212 m^3 = 2r," with"r ∈ \Z." Therefore "n ^3 \u2212 m^3" is even.